three times a number is 41 more than twice the number what is the number
Answer by Maths68(1474)

Let
Number = x
Twice of x = 2x
Three times of x = 3x
3x=2x+41
3x-2x=41
x=41
Number is 41
Twice of x = 2*41 = 82
Three times of x = 3*41=123
123=82+41
123=123